A secretary makes 2 errors per page, on average. What is the probability that on the next page she will make (a) 3 or more errors; (b) no errors?


1
Expert's answer
2021-05-07T09:42:55-0400

Average error per page "\lambda=2"


Let X denote the number of Errors.


(a) "P(X\ge3)=1-P(X<3)"


"=1-[P(X=0)+P(X=1)+P(X=2)]\\"


"=1-[\dfrac{e^{-\lambda}.\lambda^0}{0!}+\dfrac{e^{-\lambda}.\lambda^1}{1!}+\dfrac{e^{-\lambda}.\lambda^2}{2!}]"


"=1-[\dfrac{e^{-2}.2^0}{0!}+\dfrac{e^{-2}.2^1}{1!}+\dfrac{e^{-2}.2^2}{2!}]"


"=1-(0.135+0.2706+0.2706)\\=1-0.6764\\=0.3236"


(b) "P(X=0)=\dfrac{e^{-\lambda}.\lambda^0}{0!}"

"=\dfrac{e^{-2}.2^0}{0!}

 =0.135"